# Post-detection inference for sequential changepoint localization

### v2

@@ Abstract
This paper addresses a fundamental but largely unexplored challenge in sequential changepoint analysis: conducting inference following a detected change. We study the problem of localizing the changepoint using only the data observed up to a data-dependent stopping time at which a sequential detection algorithm $\mathcal{A}$ declares a change. We first construct confidence sets for the unknown changepoint when pre- and post-change distributions are assumed to be known. We then extend our framework to composite pre- and post-change scenarios. We impose no conditions on the observation space or on $\mathcal{A}$ — we only need to be able to run $\mathcal{A}$ on simulated data sequences. In summary, this work offers both theoretically sound and practically effective tools for sequential changepoint localization.

### v3

@@ Abstract
This paper addresses a fundamental but largely unexplored challenge in sequential changepoint analysis: conducting inference following a detected change. We construct confidence sets for the unknown changepoint using only the data observed up to a data-dependent stopping time at which an arbitrary sequential detection algorithm $\mathcal{A}$ declares a change. Our framework is nonparametric, makes no assumption on the composite post-change class. We extend it to the composite pre-change class under a suitable assumption. We impose no conditions on the observation space or on $\mathcal{A}$. We thus present the first general tool for sequential changepoint localization, which is theoretically sound and practically effective.